OpenStudy (anonymous):

solve for x: log(8x + 9) = 2 + log(3x − 4)

OpenStudy (anonymous):

log(8x)*log(9) = 2+log(3x)/log(4) x=2.72152

OpenStudy (kira_yamato):

\[\log \left( \frac{8x + 9}{3x - 4} \right) = 2\] \[\frac{8x + 9}{3x -4} = 100\] \[8x + 9 = 300x - 400\] \[292x = 409\] \[x = 409/292 = 1 \frac{117}{292}\]

OpenStudy (anonymous):

kira, how does 2 = 100?

OpenStudy (kira_yamato):

When you drop the log, it becomes 10^2 = 100...

OpenStudy (anonymous):

he assumed that log had base 10

OpenStudy (anonymous):

is that the case?

OpenStudy (anonymous):

thanks so much! makes sense finally lol

OpenStudy (vishweshshrimali5):

log(8x+9) = log10^2 + log(3x-4) => log(8x+9) = log (3x-4)(100) => 8x+9 = 300x - 400 => 292 x = 409 => x = 409/292

OpenStudy (vishweshshrimali5):

log(8x+9) = log10^2 + log(3x-4) => log(8x+9) = log (3x-4)(100) => 8x+9 = 300x - 400 => 292 x = 409 => x = 409/292

OpenStudy (vishweshshrimali5):

log(8x+9) = log10^2 + log(3x-4) => log(8x+9) = log (3x-4)(100) => 8x+9 = 300x - 400 => 292 x = 409 => x = 409/292